# Towards Robust Assessment of Pathological Voices via Combined Low-Level Descriptors and Foundation Model Representations

### v3

@@ Abstract
Perceptual voice quality assessment is essential for diagnosing and monitoring voice disorders by providing standardized evaluations of vocal function. Traditionally, expert raters use standard scales such as the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V) and Grade, Roughness, Breathiness, Asthenia, and Strain (GRBAS). However, these metrics are subjective and prone to inter-rater variability, motivating the need for automated, objective assessment methods. This study proposes Voice Quality Assessment Network (VOQANet), a deep learning-based framework with an attention mechanism that leverages a Speech Foundation Model (SFM) to extract high-level acoustic and prosodic information from raw speech. To enhance robustness and interpretability, we also introduce VOQANet+, which integrates low-level speech descriptors—such as jitter, shimmer, and harmonics-to-noise ratio (HNR)—with SFM embeddings into a hybrid representation. Unlike prior studies focused only on vowel-based phonation (PVQD-A subset) of the Perceptual Voice Quality Dataset (PVQD), we evaluate our models on both vowel-based and sentence-level speech (PVQD-S subset) to improve generalizability. Results show that sentence-based input outperforms vowel-based input, especially at the patient level, underscoring the value of longer utterances for capturing perceptual voice attributes. VOQANet consistently surpasses baseline methods in root mean squared error (RMSE) and Pearson correlation coefficient (PCC) across CAPE-V and GRBAS dimensions, with VOQANet+ achieving even better results. Additional experiments under noisy conditions show that VOQANet+ maintains high prediction accuracy and robustness, supporting its potential for real-world and telehealth deployment. These findings demonstrate the effectiveness of combining SFM embeddings with domain-informed acoustic features for interpretable and resilient pathological voice quality assessment.
@@ 1 Introduction
Voice disorders are common in modern society, and pathological voice quality can seriously affect an individual’s communication ability and social well-being [1], [2]. Voice disorders can be caused by a variety of diseases, including vocal fold nodules, polyps, paralysis, neurological diseases, such as Parkinson’s disease, and head and neck cancers [4]. These diseases often affect vocal characteristics, such as hoarseness, breathiness, roughness, or strain, which require perceptual examination by trained clinicians. Therefore, vocal signal analysis has become a non-invasive and accessible screening tool and is increasingly being adopted by otolaryngology and neurology clinics to help early detection and monitoring of voice-related disorders [5]. Auditory-perceptual assessment of pathological voice quality (VQA) aims to improve the diagnosis, monitoring, and treatment of voice disorders by providing an objective and standardized assessment of vocal function. It serves as a critical tool for identifying pathological voice conditions, tracking disease progression, and evaluating the effectiveness of therapeutic interventions. Traditionally, VQA relies on perceptual assessments by experienced clinicians using standardized scales, such as the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V) and the Grade, Roughness, Breathiness, Asthenia, Strain (GRBAS) [6],[7],[8]. CAPE-V has been adapted to multiple languages, including French, Turkish, European Portuguese, and Japanese, further demonstrating its clinical relevance and international applicability [9], [10],[11], [12]. It provides continuous ratings (0–100) for perceptual attributes, whereas GRBAS uses a discrete 4-point ordinal scale (0–3). For both, the larger the value, the more severe the condition. Although these perceptual ratings are widely used and provide valuable qualitative insights into voice disorders [13], they are inherently subjective and prone to inter- and intra-examiner variability. Recent studies have explored other strategies, such as crowdsourcing perceptual ratings, particularly in neurological disorders like Parkinson’s disease, to improve scalability and maintain rating validity [14]. The reliance on expert raters makes standardization difficult and increases the need for automated VQA.
@@
To address the above challenges, several machine learning (ML) and deep learning methods have been explored to automate perceptual voice assessment [15]. Traditional machine learning models, such as Random Forest (RF), Support Vector Machine (SVM), and k-Nearest Neighbors (KNN), have been used to predict CAPE-V scores based on low-level speech descriptors (LLDs), also known as handcrafted acoustic features [16],[17]. Dispite providing interpretability, they often struggle with generalizability and fail to capture the complex characteristics of speech that influence perceptual voice quality ratings. To overcome these limitations, recent studies have employed deep learning and ensemble frameworks for pathological voice classification [2], demonstrating the benefits of combining multiple modalities and learned representations for more robust voice assessment. To further enhance clinical VQA, attention-based models have also been proposed for predicting GRB scores from sustained vowel phonation, showing improved accuracy over earlier neural architecture [3]. Recent advances in Speech Foundation Models (SFM), especially large-scale pre-trained models such as WavLM [18], HuBERT [19], and Whisper [20], have performed well in extracting high-level speech representations. WavLM and HuBERT are trained using self-supervised learning (SSL), while Whisper adopts a semi-supervised paradigm, leveraging large-scale audio-text pairing data collected through weak supervision. These models provide robust and transferable features, which are well-suited for downstream tasks, including VQA. Notably, WavLM incorporates a denoising pre-training objective, which may improve the model’s robustness in noisy clinical recordings; while Whisper’s multilingual training provides a wider range of phonetic coverage, although there may be a domain mismatch with disordered speech.
@@
In this study, we propose VOQANet (Voice Quality Assessment Network), a deep learning-based framework with an attention mechanism that leverages SFM embeddings of WavLM for perceptual VQA.To further enhance clinical relevance and model robustness, we further propose VOQANet+, an extended version that combines SFM embeddings with LLDs, such as jitter, shimmer, and harmonics-to-noise ratio (HNR), which have been widely used in vocal pathology analysis for their ability to capture voice irregularities and instabilities  [21]. This combination enables VOQANet+ to benefit from both high-level learned representations and complementary low-level signal-based features. We evaluate our models on the Perceptual Voice Quality Dataset (PVQD), reporting both utterance-level and patient-level results, where the latter predictions are averaged across each speaker’s utterances. Experimental results show that VOQANet provides a strong baseline, while VOQANet+ consistently improves prediction accuracy and generalization, especially under noisy conditions.
@@ 2.1 Automated Pathological Voice Quality Assessment
Ensuring robustness and generalizability is critical for real-world applications of perceptual VQA. Traditional methods rely on machine learning models, such as RF, SVM, and KNN, which utilize LLDs, like jitter, shimmer, zero crossing rate, and HNR [17]. A lightweight feature extraction method has been proposed to leverage these classical ML models for CAPE-V prediction. However, while these models provide interpretability and domain relevance, they often struggle to generalize across datasets due to the variability of speech patterns and the sensitivity of LLDs to recording conditions [17]. While CAPE-V provides continuous ratings and GRBAS uses a discrete scale, both are susceptible to inter-rater differences [22], [23], which further motivates the development of objective, automated assessment methods.
To overcome these limitations of classical methods, recent studies have adopted deep learning-based methods, especially leveraging SFMs such as Whisper [20] and WavLM [18], which learn rich acoustic representations from large-scale raw waveforms. Among these models, WavLM includes a denoising pre-training objective, which enables the model to learn robust representations even in the presence of background noise and acoustic variations. This feature is particularly beneficial for disordered speech, which often deviates from typical acoustic patterns. On the other hand, Whisper is trained on a large-scale multilingual and multitask corpus, which makes it highly effective in a variety of automatic speech recognition (ASR) tasks. Whisper has been explored for other applications, such as improving the performance of speaker verification (SV) tasks [24]. These findings highlight the potential of SFMs in clinical applications, while also underscoring the need to tailor the representations to domain-specific characteristics. However, SFM embeddings are typically learned from general speech corpora and may not fully capture task-specific or clinically salient characteristics relevant to voice pathology. To overcome this limitation, recent studies have explored hybrid modeling strategies that combine the generalization capabilities of SFM embeddings with LLDs to improve the prediction performance of VQA systems. This has shown that such hybrid audio representations can simultaneously enhance robustness and interpretability in cognitively and physically demanding speech tasks  [25]. Based on these insights, we explore deep learning-based methods that combine SFM representations with clinically relevant feature representations.
@@ 2.2 Speech Foundation Models
Recent advances in self-supervised learning have introduced SFMs, which provide more robust and generalizable representations for downstream speech tasks. These models have attracted much attention in the speech processing community due to their ability to learn meaningful representations directly from raw audio. Models such as WavLM and HuBERT are trained using SSL, where contextual representations are learned from unlabeled data, while Whisper takes a semi-supervised approach, leveraging large-scale audio-text pairing data collected with weak supervision. Compared to traditional supervised methods, SFMs trained using SSL or semi-supervised methods can capture low-level acoustic features and high-level linguistic patterns without the need for extensive annotations.
Models pre-trained on large-scale corpora such as Librispeech  [26] have achieved excellent performance in various speech tasks such as ASR, SV, speech synthesis, and speech emotion recognition [27], [28]. In the context of VQA, SFMs have been explored for their ability to extract rich and contextualized acoustic representations that can be related to perceptual voice attributes such as breathiness, strain, and roughness. These deep representations are able to capture complex speech patterns that are usually difficult to model using only LLDs.
@@ 2.3 Hybrid Models Combining SFM and LLDs
Combining SFM embeddings with LLDs has been explored to capture data-driven and clinically interpretable speech characteristics. A hybrid approach combining data-driven features from a BYOL-derived model with LLDs extracted using openSMILE has demonstrated strong performance on speech analysis tasks [29]. Similarly, self-supervised speech models jave been shown to outperform traditional acoustic features in predicting parameters such as heart activity, highlighting the efficacy of combining SSL embeddings with LLDs [25]. While prior studies have explored hybrid designs in affective or biomedical speech tasks, few have systematically examined such approaches in the context of clinical voice assessment, particularly with respect to GRBAS and CAPE-V prediction. Although clinical findings suggest that CAPE-V ratings may correlate more strongly with objective acoustic and aerodynamic measures than GRBAS in certain populations [30], both scales remain widely used in perceptual VQA and offer complementary insights. Therefore, we include both in this study.
@@ 2.4 Evaluation Strategies in Voice Assessment Models
Assessment strategies play a critical role in evaluating the reliability and clinical applicability of voice assessment models. Most previous studies on automated VQA focus on predicting perceptual scores at the utterance level, where each audio segment is treated as an independent sample [17, 29]. While this approach enables fine-grained analysis, it may not capture broader patterns of the patient’s overall vocal characteristics, especially when there are phonetic differences between different utterances. To improve reliability in clinical settings, some studies have begun to explore patient-level assessments by aggregating predictions from multiple utterances from the same individual [25]. This strategy better reflects the real-world diagnostic process, where clinicians assess pathological voice quality based on a complete set of speech samples rather than isolated fragments.
@@ 3.1 VOQANet: SFM-Based Feature Learning
As shown in Fig. 1(a), VOQANet leverages SFM embeddings extracted by a pre-trained model to capture rich acoustic and prosodic characteristics of the input waveform. Given a waveform $\mathbf{w}$, the SFM embedding is calculated as a weighted sum of the hidden representations of all transformer layers:
$X_{S}=\sum_{\ell=0}^{L}\alpha_{\ell}\cdot h^{(\ell)}(\mathbf{w}),$ (1)
where $h^{(\ell)}$ represents the hidden state at layer $\ell$ of the pre-trained SFM, and $\alpha_{\ell}$ is a learnable scalar weight normalized by softmax. This layer-wise aggregation strategy follows prior work that showed that weighted combinations of intermediate layers outperform single-layer embeddings in speech assessment tasks [31]. This mechanism allows the model to adaptively emphasize the layers most relevant to pathological voice quality.
@@ 3.2 VOQANet+: Joint Representation Learning
As shown in Fig. 1(b), to further enhance the model’s ability to capture clinically relevant speech characteristics, VOQANet+ combines LLDs with SFM embeddings. These LLDs include Jitter, Shimmer, and HNR, which are extracted from the same waveform $\mathbf{w}$ using signal processing techniques:
$X_{A}=f_{A}(\mathbf{w}),$ (2)
@@ 3.3 Model Architecture and Training
$H=\sigma(W_{3}\cdot(\sigma(W_{2}\cdot(\sigma(W_{1}X+b_{1}))+b_{2}))+b_{3}),$ (4)
where $H$ is the latent representation, and $\sigma(\cdot)$ denotes the ReLU function. To make the network focus on the salient components of the learned features, attention-based pooling is applied after the last hidden layer to calculate a weighted summary of the feature representation. The attention module first projects $H$ into an intermediate space using a non-linear transformation, and then calculates the attention weights over the feature dimension:
$\alpha=\text{softmax}(W_{\text{attn}}\cdot\tanh(W_{h}H+b_{h})+b_{\text{attn}}).$ (5)
The attended feature vector $\mathbf{z}$ is obtained as follows,
$\mathbf{z}=\sum_{t=1}^{T}\alpha_{t}\cdot H_{t},$ (6)
where $T$ is the length of the feature vector sequence, and $H_{t}$ represents the feature vector at position $t$ in the sequence. Finally, $\mathbf{z}$ is fed into the final regression layer to predict the VQA score.

### v4

@@ Abstract
Perceptual voice quality assessment plays a vital role in diagnosing and monitoring voice disorders. Traditional methods, such as the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V) and the Grade, Roughness, Breathiness, Asthenia, and Strain (GRBAS) scales, rely on expert raters and are prone to inter-rater variability, emphasizing the need for objective solutions. This study introduces the Voice Quality Assessment Network (VOQANet), a deep learning framework that employs an attention mechanism and Speech Foundation Model (SFM) embeddings to extract high-level features. To further enhance performance, we propose VOQANet+, which integrates self-supervised SFM embeddings with low-level acoustic descriptors—namely jitter, shimmer, and harmonics-to-noise ratio (HNR). Unlike previous approaches that focus solely on vowel-based phonation (PVQD-A), our models are evaluated on both vowel-level and sentence-level speech (PVQD-S) to assess generalizability. Experimental results demonstrate that sentence-based inputs yield higher accuracy, particularly at the patient level. Overall, VOQANet consistently outperforms baseline models in terms of root mean squared error (RMSE) and Pearson correlation coefficient across CAPE-V and GRBAS dimensions, with VOQANet+ achieving even greater performance gains. Additionally, VOQANet+ maintains consistent performance under noisy conditions, suggesting enhanced robustness for real-world and telehealth applications. This work highlights the value of combining SFM embeddings with low-level features for accurate and robust pathological voice assessment.
@@ 1 Introduction
Voice disorders are common in modern society, and pathological voice quality can seriously affect an individual’s communication ability and social well-being [1], [2]. They arise from various conditions, including vocal fold nodules, polyps, paralysis, neurological diseases, such as Parkinson’s disease, and head and neck cancers [4]. These disorders affect vocal characteristics, such as hoarseness, breathiness, roughness, or strain, requiring perceptual examination by trained clinicians. Therefore, vocal signal analysis has become a widely used non-invasive screening tool in otolaryngology, neurology, and speech-language pathology clinics [5, 6]. Voice Quality Assessment (VQA) aims to improve the diagnosis, monitoring, and treatment of voice disorders by providing an objective and standardized assessment of vocal function. It serves as a critical tool for identifying pathological voice conditions, tracking disease progression, and evaluating the effectiveness of therapeutic interventions. Traditionally, VQA relies on perceptual assessments by experienced clinicians using standardized scales, such as the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V) and the Grade, Roughness, Breathiness, Asthenia, Strain (GRBAS) [7],[8],[9]. CAPE-V has been adapted to multiple languages, including French, Turkish, European Portuguese, and Japanese, further demonstrating its clinical relevance and international applicability [10], [11],[12], [13]. It provides continuous ratings (0–100) for perceptual attributes, whereas GRBAS uses a discrete 4-point ordinal scale (0–3). Recently, CAPE-V has been revised to CAPE-Vr (Consensus Auditory-Perceptual Evaluation of Voice—Revised) [14], with updated recommendations for clinical use. It has been noted that clinicians experienced challenges in rating attributes such as overall severity, strain, and pitch when using the original CAPE-V, highlighting the need for clearer definitions and more consistent administration. CAPE-Vr addresses these issues through a revised rating form, updated stimuli, and expanded categories, thereby providing more precise guidance for clinical assessment while maintaining the intent of the original protocol. For both, the larger the value, the more severe the condition. Although these perceptual ratings are widely used and provide valuable qualitative insights into voice disorders [15], they are inherently subjective and prone to inter- and intra-examiner variability. Recent studies have explored other strategies, such as crowdsourcing perceptual ratings, particularly in neurological disorders like Parkinson’s disease, to improve scalability and maintain rating validity [16]. The reliance on expert raters makes standardization difficult and increases the need for automated VQA.
@@
Machine learning (ML) and deep learning approaches have been explored to address these challenges [17]. Traditional ML models, such as Random Forest (RF), Support Vector Machine (SVM), and k-Nearest Neighbors (KNN), have been used to predict CAPE-V scores based on low-level speech descriptors (LLDs) [18],[19], but struggle to capture the complexity of pathological voices. To overcome these limitations, recent studies have employed deep learning and ensemble frameworks for pathological voice classification [2], demonstrating the benefits of combining multiple modalities and learned representations for more robust voice assessment. To further enhance clinical VQA, attention-based models have also been proposed for predicting GRB scores from sustained vowel phonation, showing improved accuracy over earlier neural architecture [3]. Recent advances in Speech Foundation Models (SFMs), including WavLM [20], HuBERT [21], and Whisper [22], have performed well in extracting high-level speech representations. WavLM and HuBERT are trained using self-supervised learning (SSL), while Whisper adopts a semi-supervised paradigm, leveraging large-scale audio-text pairing data collected through weak supervision. Beyond their success in general speech processing, HuBERT has been shown to be particularly effective in detecting pathological voices [23]. Whisper has also been explored in clinical speech-language applications, including post-stroke speech and language assessment [24]. Furthermore, WavLM has been employed for pathological voice analysis in systematic reviews and applied studies [25]. These models provide robust and transferable features, which are well-suited for downstream tasks, including VQA. WavLM’s denoising pre-training and Whisper’s multilingual coverage further motivate their use in VQA.
@@
In this study, we propose VOQANet (Voice Quality Assessment Network), a deep learning-based framework with an attention mechanism that leverages SFM embeddings for perceptual VQA. We further introduce VOQANet+, which integrates LLDs such as jitter, shimmer, and harmonics-to-noise ratio (HNR) [26] with SFM embeddings to enhance robustness and clinical interpretability. This combination enables VOQANet+ to benefit from both high-level learned representations and complementary low-level signal-based features. Both models are evaluated on the Perceptual Voice Quality Dataset (PVQD), reporting both utterance-level and patient-level results, where the latter predictions are averaged across each speaker’s utterances. Experimental results show that VOQANet provides a strong baseline, while VOQANet+ consistently improves prediction accuracy and generalization, especially under noisy conditions.
@@ 2.1 Automated Pathological Voice Quality Assessment
Ensuring robustness and generalizability is critical for real-world applications of perceptual VQA. Traditional methods rely on ML models, such as RF, SVM, and KNN using LLDs, like jitter, shimmer, zero crossing rate, and HNR [19]. A lightweight feature extraction method has been proposed to leverage these models for CAPE-V prediction. However, despite offering interpretability and domain relevance, such models often struggle to generalize across datasets due to speech variability and the sensitivity of LLDs to recording conditions [19]. While CAPE-V provides continuous ratings and GRBAS uses a discrete scale, both are susceptible to inter-rater differences [27], [28], further motivating objective and automated assessment.
To overcome these limitations, recent work has adopted deep learning-based methods, especially leveraging SFMs such as Whisper [22], and WavLM [20], which learn rich acoustic representations from large-scale raw waveforms. WavLM includes a denoising pre-training objective that improves robustness to background noise and acoustic variability. This feature is particularly beneficial for disordered speech, which often deviates from typical acoustic patterns. On the other hand, Whisper is trained on a large-scale multilingual and multitask corpus, achieves strong performance in ASR, and has been explored in tasks such as speaker verification (SV) [29]. These findings highlight the potential of SFMs in clinical applications but also underscore that embeddings learned from general speech may not fully capture clinically salient characteristics relevant to voice pathology. To overcome this limitation, hybrid approaches that combine SFM embeddings with LLDs have been explored, showing improved robustness and interpretability in demanding speech tasks [30]. Based on these insights, we explore deep learning-based methods that combine SFM representations with clinically relevant feature representations.
@@ 2.2 Speech Foundation Models
Recent advances in SSL have introduced SFMs, which provide more robust and generalizable representations for downstream speech tasks. These models have attracted much attention in the speech processing community due to their ability to learn meaningful representations directly from raw audio. Models such as WavLM and HuBERT are trained using SSL, learning contextual representations from unlabeled data, whereas Whisper adopts a semi-supervised approach leveraging large-scale audio–text pairs collected with weak supervision. Compared to traditional supervised methods, SFMs can capture low-level acoustic features and high-level linguistic patterns without the need for extensive annotations.
Models pre-trained on large-scale corpora such as Librispeech [33] have achieved excellent performance in various tasks such as ASR, SV, speech synthesis, and speech emotion recognition [34, 35]. In VQA, SFMs have been explored for their ability to extract rich, contextualized acoustic representations linked to perceptual attributes such as breathiness, strain, and roughness. These deep representations are able to capture complex speech patterns that are usually difficult to model using only LLDs.
@@ 2.3 Hybrid Models Combining SFM and LLDs
Combining SFM embeddings with LLDs has been explored as a way to capture both data-driven and clinically interpretable speech characteristics. A hybrid approach that integrates a BYOL-derived model with LLDs extracted using openSMILE has shown strong performance in speech analysis tasks [36]. Similarly, self-supervised learning (SSL)–based speech models have outperformed traditional acoustic features and physiological parameters such as heart-rate–related measures, further demonstrating the effectiveness of combining SSL embeddings with LLDs [30]. In [31] and [32], the combined features derived from SSL-based representations and LLDs achieved improved performance in both speech enhancement and speech assessment tasks, respectively, compared with using either feature type alone. In this work, we explore, for the first time, a hybrid design for clinical pathological voice assessment, targeting both GRBAS and CAPE-V prediction. Although prior clinical findings suggest that CAPE-V ratings may correlate more strongly with objective acoustic and aerodynamic measures than GRBAS in certain populations [37], both scales remain widely used in perceptual voice quality assessment and provide complementary insights. Therefore, we include both GRBAS and CAPE-V in this study.
@@ 2.4 Evaluation Strategies in Voice Assessment Models
Assessment strategies play a critical role in evaluating the reliability and clinical applicability of voice assessment models. Most prior studies on automated VQA focus on predicting perceptual scores at the utterance level, treating each audio segment as an independent sample [19, 36]. While this approach enables fine-grained analysis, it may not capture broader patterns of the patient’s overall vocal characteristics, especially when phonetic content varies across utterances. To enhance clinical reliability, recent studies have explored patient-level assessments by aggregating predictions across multiple utterances from the same speaker [30]. This strategy better reflects the real-world diagnostic process, where clinicians assess pathological voice quality based on a complete set of speech samples rather than isolated segments.
@@ 3.1 VOQANet: SFM-Based Feature Learning
As shown in Fig. 1(a), VOQANet leverages SFM embeddings extracted by a pre-trained model to capture rich acoustic and prosodic characteristics of the input waveform. Given a waveform w, the SFM embedding is calculated as a weighted sum of the hidden representations of all transformer layers:
$\boldsymbol{X}_{S}=\sum_{\ell=0}^{L}\alpha_{\ell}\cdot\mathbf{h}^{(\ell)}(\textit{{w}}),$ (1)
where $\mathbf{h}^{(\ell)}$ represents the hidden state at layer $\ell$ of the pre-trained SFM, and $\alpha_{\ell}$ is a learnable scalar weight normalized by softmax. This layer-wise aggregation strategy follows prior work that showed that weighted combinations of intermediate layers outperform single-layer embeddings in speech assessment tasks [38]. This mechanism allows the model to adaptively emphasize the layers most relevant to pathological voice quality.
@@ 3.2 VOQANet+: Joint Representation Learning
As shown in Fig. 1(b), to further enhance the model’s ability to capture clinically relevant speech characteristics, VOQANet+ combines LLDs with SFM embeddings. These LLDs include Jitter, Shimmer, and HNR, which are extracted from the same waveform w using signal processing techniques:
$\boldsymbol{X}_{A}=f_{A}(\textit{{w}}),$ (2)
@@ 3.3 Model Architecture and Training
$\textbf{H}=\sigma(\mathbf{W}_{3}\cdot(\sigma(\mathbf{W}_{2}\cdot(\sigma(\mathbf{W}_{1}\boldsymbol{X}+\mathbf{b}_{1}))+\mathbf{b}_{2}))+\mathbf{b}_{3}),$ (4)
where $\mathbf{H}$ is the latent representation, and $\sigma(\cdot)$ denotes the ReLU function. To make the network focus on the salient components of the learned features, attention-based pooling is applied after the last hidden layer to calculate a weighted summary of the feature representation. The attention module first projects $\mathbf{H}$ into an intermediate space using a non-linear transformation, and then calculates the attention weights over the feature dimension:
$\alpha=\text{softmax}(\mathbf{W}_{\text{attn}}\tanh(\mathbf{W}_{h}\mathbf{H}+\mathbf{b}_{h})+\mathbf{b}_{\text{attn}}).$ (5)
where $\mathbf{W}_{h}$ and $\mathbf{b}_{h}$ are the learnable parameters of the hidden projection that map $\mathbf{H}$ into an attention space, and $\mathbf{W}_{\text{attn}}$ and $\mathbf{b}_{\text{attn}}$ are the parameters of the output projection used to compute the attention logit. The attended feature vector z is obtained as follows,
$\textit{{z}}=\sum_{t=1}^{T}\alpha_{t}\cdot\mathbf{H}_{t},$ (6)
where $T$ is the length of the feature vector sequence, and $\mathbf{H}_{t}$ represents the feature vector at position $t$ in the sequence. Finally, z is fed into the final regression layer to predict the VQA score.